Wednesday 21st October 2009 APM Meeting tables 11:30-13:00 Present: MJI, RWA, EGS, JRL, JPE, MR, SC, STH Apologies: NAW, RGM Agenda ------ 1. Actions from last meeting 2. Comments on WFAU minutes 3. Recent and upcoming meetings 4. Data archives update - AAT, ING, WFCAM, UKIRT 5. Optical/NIR processing - HAWKI, INT WFC, MegaCam, Subaru, VST ..... 6. WFCAM update 7. VISTA update 8. AOB - grant status etc.. Minutes ------- 1. Actions from the last meeting MJI stage I software release on hold due to other pressures - ongoing <<<< EGS MJI ditto technical paper - ongoing <<<< MJI ditto utility to convert binary tables plus corrections - ongoing <<<< JRL duly wrote his ADASS talk and mailed it to the webpage STH briefly described the Palomar Transient factory. managed by Nicholas Law, it uses the 48-inch Oschin Schmidt to do an optical survey of the sky in 8 square degree snapshots and has so far produced two papers. MJI still comparing WFCAM dome flats with twilight sky flats. Low light level flats have been taken for VISTA to see how the linearity compares to WFCAM EGS has fixed the entry into the top level pages of VISTA similar to that of WFCAM on the Plone system RWA has inserted the VISTA performance specs from Sep 9 onto the VISTA technical page SC said that the Chubb engineer has not yet visited the Kavli so no progress on the high temperature monitor. Another discussion with Paul Aslin required. <<<<<<< 2. Comments on the WFAU minutes DR6 happened about a week ago. It does not include WCS but does include GPS. Edinburgh still require FITS examples so JRL will send off today - in particular that for the night of Sept 27. Also the form of the FITs headers in tile files has been decided. STH said that these adhered to the FITS standard whilst the WFCAM files did not. FITS information will be placed in the primary header of the catalogue. 3. Recent and upcoming meetings There was an IOT on Oct 6 in which science verification and scheduling was discussed. SV started on Oct 15 and runs until the end of the month. ESO wanted to produce scientific quality data at the telescope but JRL thought this was not going to happen. We'll process data here and then visit ESO in November to discuss data processing directly. JRL reported that ADASS was successful, particularly in discussions after his talk when he was subjected to many questions about VISTA. MJI and NAW went to a meeting on Euclid in Durham. Although this is a dark energy mission, three of its cameras were of interest, particularly a near-IR one which would scan 20000 square degrees of high latitude sky down to 24th mag in 5 minutes (JPE held head in hands here, but cheered up when he heard the launch is in 2018). MJI suggested we keep an interest in this and PLATO at a low level. STH said there was a PLATO meeting in early December involving an STFC bid. 4. Data archives AAT, ING and WFCAM are up-to-date but output from UKIRT is at a low level at the moment. We are still awaiting 2 to 3 dozen nights to complete the data archive for UKIRT. JRL promises to do this by Christmas. 5, Optical /NIR processing All of the current optical processing in the group is for science research projects. Volumes from VST were expected by Nov 2010. 6. WFCAM update WFCAM has been taken off standby and is now back on the telescope. MJI has been checking the data and has found an anomaly on detector 4 due to thermal cycling and more pick-up on detector 3 which does not appear to be correlated with the other quadrants. MJI said that the Rice compression programme for 16 bit data had been updated for CFITSIO and that we should warn ROE about this. The new DS9 will be able to read this. MJI looked forward to the next six-monthly UKIRT Board Meeting and needs to report on these matters as well as giving a grant update. Reprocessing is on hold because of VISTA but there is not much to do - GPS nebulosity and hot stack UPS. DR6 has been fixed and released. EGS has produced some nice statistics on UKIDSS for Steve Warren. UKIRT processing will start again next week and new flats will be needed after the standdown. 7. VISTA update JPE brought his own disks of VISTA data containing nights from Sept 27 to Oct 8 to IoA on Oct 10 where it was read in immediately by MJI, although not without problems. A copy of all this data appeared on the usual ESO delivery five days later. This is all PR stuff and contains images of galactic centre, the Fornax cluster and IC1365. This data is still being processed and several bugs found but some PR images are with Richard Hook. JPE thought the next ESO disk would be mostly empty and not contain any SV or publicity data. We are not using the ESO pipeline for this. JRL pointed out some problems with flat fields due to the close proximity of a 5th mag star whose diffraction spike was occasionally in the field of view when the telescope was jittered, so care needs to be taken when co-adding flats. Another problem arises when taking short exposures. Steven Beard has a routine for calculating exposure times for given count rates but it does not allow for readout time (1 second) so at short exposures it is easy to overexpose the frame. The rule of thumb is that if exposure time is less than 5 seconds then the sky is too bright, so do not use short exposures or those already obtained which are affected. It was suggested that flats were done in the half hour before sunrise but JPE confirmed this was not happening at present. This prompted the comment that saturating WFCAM with bright skylight at dusk produces a remanence which affects later exposures, and it is not yet known whether this feature affects VISTA. The dome lights have been changed and the new transformers mean that the lights are left on for an hour before use. MJI stressed the urgency of using them for non-linearity sequences and JRL said he had already enquired about this. The recent attempts to speed up delivery of data to CASU prompted suggestions that this should be repeated for SV data. JRL said we need guidelines for offset sky observations. No one has yet tried Steven Beard's offset sky template amd JPE said that ESO would probably change it after the Dry Runs. It seems that OBs stop when the AG drops out but JPE said that it was possible to restart a pawprint if this happens in the middle of the procedure. MJI said that he would build in safeguards to the preview script. <<<<<<<<< The PR pictures yielded some nice images of the Fornax Cluster and a deadline for the release of these was understood to be Nov 4 at present although JPE said that this may slip by one or two weeks. He does not know when VISTA will be accepted by ESO. JPE described the ESOcasts which are propagated around the site and showed some of the publicity images. MJI said he was keen to get the PR exercise out of the way before the SV data comes in. MJI talked about mosaicing, starting with the different mosaicing and tiling software. The default program is Montage but this uses 64-bit floating point and non-standard FITS headers which are drawbacks. We should produce our own prototype running 32-bit with sensible headers and will soon be able to send data reduced in this fashion to ROE. He explained the differences between arrays such as HAWK-I, VISTA and WFCAM. The current idea is that we should generate catalogues and then go back to the pawprints to update the photometry. STH said that he had done tests which showed that the differences between pawprints and tiles was about 5%. STH is checking colour equations and throughput and agrees with the data on the webpage. He confirms that the low value for WFCAM (16%) is correct and this is due to the much lower QE than the VISTA detectors. STH is redoing zero points at the moment. JPE said that he was worried about possible time-dependent astrometric distortion in the focal plane asked for it to be monited by the QC pipeline. MJI said that he would occasionally look at distortion patterns by using stacking. <<<<<<<<< Do objects move around in the focal plane when different narrow band filters are used. Mike said it was nothing we could deal with at present but it is not simple illumination error. JPE explained that there were 2 sets of eight ND filters in the cryostat which had never been exposed to light and Tim de Zeeuw had allocated 3 nights for them to be assessed. JPE said that he himself had 6 nights to assess the Y and Z filters. JPE will give MJI the filter data for the NB filters and NB118. <<<<<<<< There is still disagreement with ESO and CASU about the use of the ETC. MJI said that ESO were using the wrong sky background. The old version by PSB has been found and removed. MJI said that the stability of the filter wheel was still being monitored - the last problem occurred on Sept 29 and has not been seen since. JPE said he meant to re-datum each night but did not always do so. JPE said that microstepping was required when the seeing was ? better/worse than ? MJI disagreed that this needed to be done as we have to interpolate anyway and this can be done on a fine grid. STH said that he would look at image quality again when more decent data is available. He expected that the SMC data on Oct 5-7 would be suitable and would soon be accessible. There followed some discussion about the PSF profile and its dependence on what type of profile is assumed. <<<<<<<< STH was keep to see trail survey OBs before they are sorted and suggested that we should have access to them at the same time as ESO otherwise it is too late. JPE suggested that ESO would not do this but that the PIs should, however he would ask ESO. STH and MJI wanted to do science with the SV data and suggested they should be on the SV panel. 8. Any other business All of the GAIA hardware is now running and MJI said that 2 to 3 out of the 4 A/C units were running at any one time. Both JRL and EGS have had problems with the power supply to their PCs recently and suffer unduly from mains outages. An anti-surge unit would be installed in their office in due course. <<<<<<<<< Non-cluster machine backups had been done (i.e. apm 25, 26 28 etc...) MJI expressed the hope that we could now replace PSB on the VISTA side and expected to progress this soon with closing date by the end of November, interviews in early January for a start on Apr 1. Continuing actions ------------------ EGS add catalogue software and binary table conversion software to MJI software release pages MJI finish off technical paper updates MJI write utility to convert binary tables with all corrections New actions ----------- JRL to prod Andy Adamson into releasing the last few nights needed to complete the UKIRT archive. MJI to tell ROE about the new Rice compression routine for CFITSIO. JRL to find out when changed caused by filter wheel occurred. EGS to monitor JRLs sky frames. JRL+MJI investigate squishing VISTA images at the end of processing a la Pence.